# Irreducible finite-index representations of automorphism groups of trees

### v1

@@ 5 Preserved sesquilinear forms
$\overline{\alpha}\left(1-|\alpha|^{2}\right)B(h,h)=(1-\overline{\alpha}^{2})\alpha B(h,h).$
We can divide by $B(h,h)$ on the right as it is nonzero. By developing the two expressions we get
$\overline{\alpha}-\overline{\alpha}|\alpha|^{2}=\alpha-\overline{\alpha}|\alpha|^{2}$
and therefore $\overline{\alpha}=\alpha$, meaning that $\alpha\in\mathbb{R}$.
@@
Every spherical irreducible complex or quaternionic representation of $G$ preserving a finite-index nondegenerate sesquilinear form is obtained by tensoring with $\mathbb{C}$ or $\mathbb{H}$ a spherical irreducible real representation of $G$ preserving a finite-index nondegenerate bilinear form.
Therefore, from now on, we may assume $\mathbb{K}=\mathbb{R}$, and we are left to show that no nondegenerate bilinear form $B$ of finite index >1>1> 1 can be preserved by a spherical representation of $G$.
The strategy will be to present, for every $\alpha\in\mathbb{R}\setminus\{\pm 1\}$, an explicit bilinear form of index 0 or 111 that is preserved by the corresponding representation of $G$. This will exclude, by Proposition 3.3, that the representation can preserve a bilinear form of index >1>1> 1, proving the theorem. Furthermore, this will provide a classification of bilinear forms that can be preserved by a spherical irreducible representation. As we have said, from now on we work with real representations, but the same results will hold for complex and quaternionic representations.
@@
This bilinear form has index 0 if $|\alpha|<1$ and index 111 if $|\alpha|>1$.
@@
Let $g_{0}\in G$. To show that $g_{0}$ can be generated with $g$ and $K$, let $w:=g_{0}^{-1}(o)$. If we find $g_{1}$ generated with $g$ and $K$ such that $g_{1}^{-1}(o)=w$, we will have $g_{0}g_{1}^{-1}(o)=o$ and therefore $g_{0}g_{1}^{-1}\in K$, concluding.
To find such a $g_{1}$, let $m:=d(o,w)$. Observe that by composing $g$ with some element of $K$ we can get an element $g^{\prime}$ that sends $o$ to $v$ and $v$ to a vertex in $S_{2}$. Then the element $(g^{\prime})^{m}$ will send $o$ to a vertex in $S_{m}$, and by composing with an element of $K$ we can get $g^{\prime\prime}$ such that $g^{\prime\prime}(o)=w$. Now, it suffices to take $g_{1}=(g^{\prime\prime})^{-1}$.
@@ 6 The non-transitive case
The tree must be homogeneous or semi-homogeneous. We may assume that it is semi-homogeneous of degrees $r,s\geq 3$: the homogeneous case will be the special case $r=s$. We may also assume, without loss of generality, that the base vertex $o$ has degree $r$.
@@
Proposition 3.2 implies that $(\pi,H)$ is equivalent to a subrepresentation of $(\Pi,C)$, where $C$ is the space of functions on $W:=\{v\in V:d(o,v)\text{ even}\}$ that are left-invariant under some open compact subgroup of $G$, and $\Pi$ is the left action of $G$ on this space.
@@
This implies, similarly to the transitive case with the Laplacian, that for every $\alpha\in\mathbb{K}$ the space of radial functions that satisfy $L_{2}f=\alpha f$ is 111-dimensional.
Analogously to the transitive case, we can use the Gel’fand pair property of $(G,K)$ to show that there exists a function $f\in H$ such that the space of $K$-invariant functions in $H$ is the space of scalar multiples of $f$, and there exists $\alpha\in\mathbb{K}$ such that $L_{2}f=\alpha f$.
@@
If $\alpha\neq 1$ and $\alpha\neq-\frac{1}{s-1}$, then the natural analog of Theorem 4.6 holds: the representation $(\pi,H)$ is equivalent to the left action of $G$ on the space $\{h\in C:L_{2}h=\alpha h\}$, which is spanned by the $G$-translates of $f$, and conversely for every $\alpha\in\mathbb{K}$, $\alpha\neq 1$, $\alpha\neq-\frac{1}{s-1}$, the construction gives an irreducible spherical representation of $G$.
@@
We can repeat the proof of Theorem 4.6. The only part that needs to be changed (and that will lead to the exception for $\alpha=-\frac{1}{s-1}$) is when we show that the space $\{h\in C:L_{2}h=\alpha h\}$ is generated by the $G$-translates of the radial function $f$. To show this in our case, we fix a function $h$ in the space and try to build it with scalar multiples of translates of $f$.
We start with $F=h(o)f(0)^{-1}f$, which agrees with $h$ at $o$. Now we wish to make it agree with $h$ on $S_{2}$ as well. For every $w\in S_{2}$, let $g_{w}\in G$ send $o$ to $w$. We add to $F$ the function
$\sum_{v\in N(o)}{\sum_{w\in N(v)}{(h(w)-F(w))(f(0)-f(2))^{-1}\pi(g_{w})f}}.$
The sum of the coefficients is zero, which guarantees that $F$ is unchanged at $o$. For every $v\in N(o)$, let $R(v):=\sum_{w\in N(v)}{(h(w)-F(w))}$. The condition $L_{2}h=\alpha h$, together with $L_{2}F=\alpha F$, implies that $\sum_{v\in N(o)}{R(v)}=0$.
@@
All this adds up to $h(w)-F(w)+R(v)(f(0)-f(2))^{-1}(f(2)-f(4))$.
To remove the extra term $T(v):=R(v)(f(0)-f(2))^{-1}(f(2)-f(4))$, we subtract from $F$ the function
$\sum_{v\in N(o)}{T(v)(f(0)+(s-2)f(2)-(s-1)f(4))^{-1}}\sum_{w\in N(v)}{\pi(g_{w})f}.$
@@
All this adds up to $-T(v)$, so now the function $F$ agrees with $h$ on $S_{2}$ as well as at $o$.
@@
Therefore, the theorem is proved in the non-transitive case whenever $f(0)+(s-2)f(2)-(s-1)f(4)\neq 0$. As $f(2)=\alpha f(0)$ and
@@
$=\frac{r(s-1)\alpha^{2}-(s-2)\alpha-1}{(r-1)(s-1)}f(0),$
@@
We are left with the case $\alpha=-\frac{1}{s-1}$. In that case, the function $f$ satisfies $f(2n)=\alpha^{n}f(0)$ for all $n\in\mathbb{N}$. This is checked directly for $n=0,1$ and by induction for $n\geq 2$: we have
@@
If $\alpha=-\frac{1}{s-1}$, the space $H$ is the subspace of $\{h\in C:L_{2}h=\alpha h\}$ consisting of functions with sum zero on the set of neighbors of any vertex in $V\setminus W$ (i. e. any vertex at odd distance from $o$).
@@
[…]
We only need to show that the $G$-translates of $f$ span our space. We can follow the exact same proof of Theorem 6.2, up to the point where we are left with the term $T(v)=R(v)(f(0)-f(2))^{-1}(f(2)-f(4))$. We cannot eliminate this term with the procedure that follows in the proof, because it would involve multiplying functions by the inverse of zero. However, we can show that in our case there is nothing to eliminate, as we have $R(v)=0$.
Recall that, in fact, $R(v)$ was the sum of $h(w)-F(w)$ where $w$ varies in the set $N(v)$. If we let $\overline{v}$ be the unique neighbor of $v$ which is not in $N(v)$ (i. e. the one on the path to $o$), as the sum of $h-F$ over $N(v)\cup\{\overline{v}\}$ is zero we have $R(v)=F(\overline{v})-h(\overline{v})$. The functions $F$ and $h$ already agree at $\overline{v}$, therefore $R(v)=0$.
@@
We divide $H=\{h\in C:L_{2}h=\alpha h\}$ into a direct sum of $K$-invariant subspaces: $H=\bigoplus_{n\in\mathbb{N}}{H_{n}}$. Again, $H_{0}$ is 111-dimensional (spanned by a radial $f$). Then, for $n\geq 1$ we define $\widetilde{H}_{n}$ as the subspace of $H$ consisting of functions that are invariant under $G(B_{n}(o))$. In this way, a function in $\widetilde{H}_{n}$ will be determined by its values in $S_{2k}$ if $n\in\{2k-1,2k\}$, and in the case $n=2k-1$ the value will have to be constant on $N(w)$ for every $w\in S_{2k-1}$. We can define, for every $k\geq 1$:
@@
Like in the transitive case, we prove that $H_{1}$ is irreducible and that for every $n\geq 2$ each $K$-irreducible subspace of $H_{n}$ has dimension at least $|S_{n-1}|$. The proof is exactly the same if $n$ is even, while if $n=2k-1$ is odd, we consider, in place of functions on $S_{2k-1}$, functions on $S_{2k}$ that are constant on $N(w)$ for every $w\in S_{2k-1}$ (as is the case for functions in $H_{2k-1}$).
@@
If the representation preserves the finite-index sesquilinear form $B$ (which we assume to be antilinear in the first variable and linear in the second), then $\alpha\in\mathbb{R}$.
@@
If we consider $g\in G$ that sends $o$ to $x$ and $x$ to $o$, we have $\pi(g)f=\alpha f+(1-\alpha)h$ and $\pi(g)h=(1+\alpha)f-\alpha h$. The former implies
$|1-\alpha|^{2}B(h,h)=\left(1-|\alpha|^{2}\right)B(f,f).$
@@
Since $B(f,h)=0$, it must be $B(\pi(g)f,\pi(g)h)=0$. This means that $B(\alpha f+(1-\alpha)h,(1+\alpha)f-\alpha h)=0$, and therefore
$\overline{\alpha}(1+\alpha)B(f,f)-(1-\overline{\alpha})\alpha B(h,h)=0.$
Multiplying by $1-\alpha$ on the left and using the equality above, we get
$(1-\alpha)\overline{\alpha}(1+\alpha)B(f,f)=|1-\alpha|^{2}\alpha B(h,h)=\alpha\left(1-|\alpha|^{2}\right)B(f,f).$
We can divide by $B(f,f)$ on the right, since it is nonzero. By developing the two expressions, we get
$\overline{\alpha}-\alpha|\alpha|^{2}=\alpha-\alpha|\alpha|^{2}$
and therefore $\overline{\alpha}=\alpha$, which means that $\alpha\in\mathbb{R}$.
@@
Like in the transitive case, we finish the proof of Theorem 2.3 by presenting, for every $\alpha\in\mathbb{R}\setminus\{1\}$, an explicit bilinear form preserved under $G$. In the non-transitive case, the bilinear form will be of index 0, 111 or $\infty$: in the first two cases, Proposition 3.3 applies directly and excludes that $G$ can preserve a form of index >1>1> 1, while in the third case we can argue by contradiction that if $G$ preserved a form of finite index $p$ it could not preserve (again by Proposition 3.3) a form of index $\infty$.
@@
This bilinear form has index 0 if $-\frac{1}{s-1}<\alpha<1$, index 111 if $\alpha>1$, and index $\infty$ if $\alpha<-\frac{1}{s-1}$.
We need to check that this bilinear form is preserved under $G$. Let $v\in N(o)$ and let $w\in N(v)$. Consider $g\in G$ that sends $o$ to $w$ and $w$ to $o$. The proof of Lemma 5.6 can be easily adapted to show that $g$ and $K$ generate $G$. As the bilinear form $Q$ is clearly preserved under $K$, it suffices to show that it is preserved under $g$.
@@
Now, we need to show that $Q(h,j)=0$. We may assume (up to multiplying by a scalar) that $j(o)=1$, which implies $j(w)=-1$. Let $a$ be the value of $h$ at $o$ (or $w$) and let $b$ be its value at any $x\in D\setminus\{o,w\}$. The function $h$ is the sum of the following:
@@
  • a function in $H_{1}$ with value $\left(\frac{1}{s-1}-\alpha\right)a+\frac{s-2}{s-1}b$ at any $x\in N(v)$ and value $-\frac{1}{r-1}\left(\left(\frac{1}{s-1}-\alpha\right)a+\frac{s-2}{s-1}b\right)$ at any $x\in S_{2}\setminus N(v)$,
  • a function in $H_{2}$ supported on $N(v)$ with value $\frac{s-2}{s-1}(a-b)$ at $w$ and value $-\frac{1}{s-1}(a-b)$ at any $x\in N(v)\setminus\{w\}$.
@@
  • a function in $H_{0}$ with value 111 at $o$,
  • a function in $H_{1}$ with value $-\frac{1}{s-1}-\alpha$ at any $x\in N(v)$ and value $\frac{1}{r-1}\left(\frac{1}{s-1}+\alpha\right)$ at any $x\in S_{2}\setminus N(v)$,
  • a function in $H_{2}$ supported on $N(v)$ with value $-\frac{s-2}{s-1}$ at $w$ and $\frac{1}{s-1}$ at any $x\in N(v)\setminus\{w\}$.
@@
Therefore, we have shown that $g$ preserves $Q$ on $H_{D}$. We now show that $g$ preserves $Q$ on $H_{D}^{+}$, in a way similar to the end of the proof of Proposition 5.5.
@@
For every $n\geq 1$ we have $\pi(g)q_{n}^{o}\in H_{n+2}^{w}$, for every $n\geq 3$ we have $\pi(g)q_{n}^{x}\in H_{n}^{x}$, and for every $n\geq 3$ we have $\pi(g)q_{n}^{w}\in H_{n-2}^{o}$. From the definition of the bilinear form $Q$ it is then immediate to see that the norm of each $q_{n}^{o}$, $q_{n}^{x}$ and $q_{n}^{w}$ is preserved, and as all addends are orthogonal, it follows that the norm of $q$ is preserved. This completes the proof of Theorem 2.3 in the non-transitive case, if we assume $\alpha\neq-\frac{1}{s-1}$.
@@
[…]
The proof that $g$ preserves the three subspaces, that it preserves $Q$ on each of them, and that $H_{D}^{(3)}$ is orthogonal to both of the others works exactly as in the case $\alpha\neq-\frac{1}{s-1}$. The only point to change is where we check that $H_{D}^{(1)}$ and $H_{D}^{(2)}$ are orthogonal.
[…]

### v2

@@ 5 Preserved sesquilinear forms
$B(h,h)\left(1-|\alpha|^{2}\right)\overline{\alpha}=B(h,h)(1-\overline{\alpha}^{2})\alpha.$
We can divide by $B(h,h)$ on the left as it is nonzero. By developing the two expressions we get
$\overline{\alpha}-|\alpha|^{2}\overline{\alpha}=\alpha-\overline{\alpha}|\alpha|^{2}$
and, therefore, $\overline{\alpha}=\alpha$, meaning that $\alpha\in\mathbb{R}$.
@@
Every spherical irreducible complex or quaternionic representation of $G$ preserving a nondegenerate sesquilinear form is obtained by tensoring with $\mathbb{C}$ or $\mathbb{H}$ a spherical irreducible real representation of $G$ preserving a nondegenerate bilinear form.
Therefore, from now on, we may assume $\mathbb{K}=\mathbb{R}$, and we are left to show that no nondegenerate bilinear form $B$ of finite index >1>1 can be preserved by a spherical representation of $G$.
The strategy will be to present, for every $\alpha\in\mathbb{R}\setminus\{\pm 1\}$, an explicit nondegenerate bilinear form of index 0 or $1$ that is preserved by the corresponding representation of $G$. This will exclude, by Proposition 3.3, that the representation can preserve a nondegenerate bilinear form of index >1>1, proving the theorem. Furthermore, this will provide a classification of nondegenerate bilinear forms that can be preserved by a spherical irreducible representation. As we have said, from now on we work with real representations, but the same results will hold for complex and quaternionic representations.
@@
This bilinear form is nondegenerate with index 0 if $|\alpha|<1$ and index $1$ if $|\alpha|>1$.
@@
Let $g_{0}\in G$. To show that $g_{0}$ can be generated with $g$ and $K$, let $w:=g_{0}(o)$. If we find $g_{1}$ generated with $g$ and $K$ such that $g_{1}(o)=w$, we will have $g_{1}^{-1}g_{0}(o)=o$ and therefore $g_{1}^{-1}g_{0}\in K$, concluding.
To find such a $g_{1}$, let $m:=d(o,w)$. Observe that by composing $g$ with some element of $K$ we can get an element $g^{\prime}$ that sends $o$ to $v$ and $v$ to a vertex in $S_{2}$. Then the element $(g^{\prime})^{m}$ will send $o$ to a vertex in $S_{m}$, and by composing with an element of $K$ we can get $g_{1}$ such that $g_{1}(o)=w$.
@@ 6 The non-transitive case
The tree must be regular or biregular. We may assume that it is biregular of degrees $r,s\geq 3$: the regular case will be the special case $r=s$. We may also assume, without loss of generality, that the base vertex $o$ has degree $r$.
@@
Proposition 3.2 implies that $(\pi,H)$ is equivalent to a subrepresentation of $(\Pi,C)$, where $C$ is the space of functions on $W:=\{v\in V:d(o,v)\text{ even}\}$ that are left-invariant under some compact open subgroup of $G$, and $\Pi$ is the left action of $G$ on this space.
@@
Analogously to the transitive case, we can use the Gel’fand pair property of $(G,K)$ to show that there exists a radial function $f\in H$ with $f(0)=1$ such that the space of $K$-invariant functions in $H$ is the space of scalar multiples of $f$, and there exists $\alpha\in\mathbb{K}$ such that $L_{2}f=\alpha\cdot f$.
@@
If $\alpha\neq 1$ and $\alpha\neq-\frac{1}{s-1}$, then the natural analog of Theorem 4.6 holds: the representation $(\pi,H)$ is equivalent to the left action of $G$ on the space $\{h\in C:L_{2}h=\alpha\cdot h\}$, which is spanned by the $G$-translates of $f$, and conversely for every $\alpha\in\mathbb{K}$, $\alpha\neq 1$, $\alpha\neq-\frac{1}{s-1}$, the construction gives an irreducible spherical representation of $G$.
@@
We can repeat the proof of Theorem 4.6. The only part that needs to be changed (and that will lead to the exception for $\alpha=-\frac{1}{s-1}$) is when we show that the space $\{h\in C:L_{2}h=\alpha\cdot h\}$ is generated by the $G$-translates of the radial function $f$. To show this in our case, we fix a function $h$ in the space and try to build it with scalar multiples of translates of $f$.
We start with $F=f\cdot h(o)$, which agrees with $h$ at $o$. Now we wish to make it agree with $h$ on $S_{2}$ as well. For every $w\in S_{2}$, let $g_{w}\in G$ send $o$ to $w$. We add to $F$ the function
$\sum_{v\in N(o)}{\sum_{w\in N(v)}{\pi(g_{w})f\cdot(f(0)-f(2))^{-1}(h(w)-F(w))}}.$
The sum of the coefficients is zero, which guarantees that $F$ is unchanged at $o$. For every $v\in N(o)$, let $R(v):=\sum_{w\in N(v)}{(h(w)-F(w))}$. The condition $L_{2}h=\alpha\cdot h$, together with $L_{2}F=\alpha\cdot F$, implies that $\sum_{v\in N(o)}{R(v)}=0$.
@@
All this adds up to $h(w)-F(w)+(f(2)-f(4))(f(0)-f(2))^{-1}R(v)$.
To remove the extra term $T(v):=(f(2)-f(4))(f(0)-f(2))^{-1}R(v)$, we subtract from $F$ the function
$\sum_{v\in N(o)}{\sum_{w\in N(v)}{\pi(g_{w})f\cdot(f(0)+(s-2)f(2)-(s-1)f(4))^{-1}T(v)}}.$
@@
All this adds up to T(v)-T(v), so now the function $F$ agrees with $h$ on $S_{2}$ as well as at $o$.
@@
Therefore, the theorem is proved in the non-transitive case whenever $f(0)+(s-2)f(2)-(s-1)f(4)\neq 0$. As $f(2)=\alpha$ and
@@
$=\frac{r(s-1)\alpha^{2}-(s-2)\alpha-1}{(r-1)(s-1)},$
@@
We are left with the case $\alpha=-\frac{1}{s-1}$. In that case, the function $f$ satisfies $f(2n)=\alpha^{n}$ for all $n\in\mathbb{N}$. This is checked directly for $n=0,1$ and by induction for $n\geq 2$: we have
@@
If $\alpha=-\frac{1}{s-1}$, the space $H$ is the subspace of $\{h\in C:L_{2}h=\alpha\cdot h\}$ consisting of functions with sum zero on the set of neighbors of any vertex in $V\setminus W$ (i. e. any vertex at odd distance from $o$).
@@
As $f(2n)=\alpha^{n}$ for all $n$, it is easy to see that $f$ is in the space described in the statement: if a vertex is at distance 2k12k-1 from $o$, the function will have value $\alpha^{k-1}$ at one neighbor and $\alpha^{k}$ at all other neighbors, adding up to $\alpha^{k-1}(1+(s-1)\alpha)=0$. It is also easy to see that the space is preserved under the action of $G$.
We only need to show that the $G$-translates of $f$ span our space. We can follow the exact same proof of Theorem 6.2, up to the point where we are left with the term $T(v)=(f(2)-f(4))(f(0)-f(2))^{-1}R(v)$. We cannot eliminate this term with the procedure that follows in the proof, because it would involve multiplying functions by the inverse of zero. However, we can show that in our case there is nothing to eliminate, as we have $R(v)=0$.
Recall that, in fact, $R(v)$ was the sum of h(w)F(w)h(w)-F(w) where $w$ varies in the set $N(v)$. If we let $\overline{v}$ be the unique neighbor of $v$ which is not in $N(v)$ (i. e. the one on the path to $o$), since the sum of hFh-F over $N(v)\cup\{\overline{v}\}$ is zero we have $R(v)=F(\overline{v})-h(\overline{v})$. The functions $F$ and $h$ already agree at $\overline{v}$, therefore $R(v)=0$.
@@
We split $H=\{h\in C:L_{2}h=\alpha\cdot h\}$ into a direct sum of $K$-invariant subspaces: $H=\bigoplus_{n\in\mathbb{N}}{H_{n}}$. Again, $H_{0}$ is $1$-dimensional (spanned by a radial $f$). Then, for $n\geq 1$ we define $\widetilde{H}_{n}$ as the subspace of $H$ consisting of functions that are invariant under $G(B_{n}(o))$. In this way, a function in $\widetilde{H}_{n}$ will be determined by its values in $S_{2k}$ if $n\in\{2k-1,2k\}$, and in the case n=2k1n=2k-1 the value will have to be constant on $N(w)$ for every $w\in S_{2k-1}$. We can define, for every $k\geq 1$:
@@
Like in the transitive case, we prove that $H_{1}$ is irreducible and that for every $n\geq 2$ each $K$-irreducible subspace of $H_{n}$ has dimension at least $|S_{n-1}|$. The proof is exactly the same if $n$ is even, while if n=2k1n=2k-1 is odd, we consider, in place of functions on $S_{2k-1}$, functions on $S_{2k}$ that are constant on $N(w)$ for every $w\in S_{2k-1}$ (as is the case for functions in $H_{2k-1}$).
@@
If the representation preserves the nondegenerate sesquilinear form $B$ (which we assume to be antilinear in the first variable and linear in the second), then $\alpha\in\mathbb{R}$.
@@
If we consider $g\in G$ that sends $o$ to $x$ and $x$ to $o$, we have $\pi(g)f=f\cdot\alpha+h\cdot(1-\alpha)$ and $\pi(g)h=f\cdot(1+\alpha)-h\cdot\alpha$. The former implies
$B(h,h)|1-\alpha|^{2}=B(f,f)\left(1-|\alpha|^{2}\right).$
@@
Since $B(f,h)=0$, it must be $B(\pi(g)f,\pi(g)h)=0$. This means that $B(f\cdot\alpha+h\cdot(1-\alpha),f\cdot(1+\alpha)-h\cdot\alpha)=0$, and therefore
$B(f,f)\overline{\alpha}(1+\alpha)-B(h,h)(1-\overline{\alpha})\alpha=0.$
Multiplying by $1-\alpha$ on the left and using the equality above, recalling that $B(f,f)$ and $B(h,h)$ are in $\mathbb{R}$, we get
$B(f,f)(1-\alpha)\overline{\alpha}(1+\alpha)=B(h,h)|1-\alpha|^{2}\alpha=B(f,f)\left(1-|\alpha|^{2}\right)\alpha.$
We can divide by $B(f,f)$ since it is nonzero. Developing the two expressions, we get
$\overline{\alpha}-|\alpha|^{2}\alpha=\alpha-|\alpha|^{2}\alpha$
and therefore $\overline{\alpha}=\alpha$, which means $\alpha\in\mathbb{R}$.
@@
Like in the transitive case, we finish the proof of Theorem 1.2 by presenting, for every $\alpha\in\mathbb{R}\setminus\{1\}$, an explicit nondegenerate bilinear form preserved under $G$. In the non-transitive case, the bilinear form will be of index 0, 11 or $\infty$: in the first two cases, Proposition 3.3 applies directly and excludes that $G$ can preserve a form of index >1>1, while in the third case we can argue by contradiction that if $G$ preserved a form of finite index $p$ it could not preserve (again by Proposition 3.3) a form of index $\infty$.
@@
This bilinear form is nondegenerate with index 0 if $\frac{1}{s-1}<\alpha<1$, index $1$ if $\alpha>1$, and index $\infty$ if $\alpha<-\frac{1}{s-1}$.
We need to check that this bilinear form is preserved under $G$. Let $v\in N(o)$ and let $w\in N(v)$. Consider $g\in G$ which sends $o$ to $w$ and $w$ to $o$. The proof of Lemma 5.6 can be easily adapted to show that $g$ and $K$ generate $G$. As the bilinear form $Q$ is clearly preserved under $K$, it suffices to show that it is preserved under $g$.
@@
Now, we need to show that $Q(h,j)=0$. We may assume (up to multiplying by a scalar) that $j(o)=1$, which implies j(w)=1j(w)=-1. Let $a$ be the value of $h$ at $o$ (or $w$) and let $b$ be its value at any $x\in D\setminus\{o,w\}$. The function $h$ is the sum of the following:
@@
  • a function in $H_{1}$ with value $\left(\frac{1}{s-1}-\alpha\right)a+\frac{s-2}{s-1}b$ at any $x\in N(v)$ and value $\frac{1}{r-1}\left(\left(\frac{1}{s-1}-\alpha\right)a+\frac{s-2}{s-1}b\right)$ at any $x\in S_{2}\setminus N(v)$,
  • a function in $H_{2}$ supported on $N(v)$ with value $\frac{s-2}{s-1}(a-b)$ at $w$ and value $\frac{1}{s-1}(a-b)$ at any $x\in N(v)\setminus\{w\}$.
@@
  • a function in $H_{0}$ with value $1$ at $o$,
  • a function in $H_{1}$ with value $\frac{1}{s-1}-\alpha$ at any $x\in N(v)$ and value $\frac{1}{r-1}\left(\frac{1}{s-1}+\alpha\right)$ at any $x\in S_{2}\setminus N(v)$,
  • a function in $H_{2}$ supported on $N(v)$ with value $\frac{s-2}{s-1}$ at $w$ and $\frac{1}{s-1}$ at any $x\in N(v)\setminus\{w\}$.
@@
Therefore, we have shown that $g$ preserves $Q$ on $H_{D}$. We now show that $g$ preserves $Q$ on $H_{D}^{\perp}$, in a way similar to the end of the proof of Proposition 5.5.
@@
For every $n\geq 1$ we have $\pi(g)q_{n}^{o}\in H_{n+2}^{w}$, for every $n\geq 3$ we have $\pi(g)q_{n}^{x}\in H_{n}^{x}$, and for every $n\geq 3$ we have $\pi(g)q_{n}^{w}\in H_{n-2}^{o}$. From the definition of the bilinear form $Q$ it is then immediate to see that the norm of each $q_{n}^{o}$, $q_{n}^{x}$ and $q_{n}^{w}$ is preserved, and as all the addends are orthogonal, it follows that the norm of $q$ is preserved. This completes the proof of Theorem 1.2 in the non-transitive case if we assume $\alpha\neq-\frac{1}{s-1}$.
@@
We can split $H_{D}$ into the same direct sum $H_{D}^{(1)}\oplus H_{D}^{(2)}\oplus H_{D}^{(3)}$. The only difference is that, while in the case $\alpha\neq-\frac{1}{s-1}$ the space $H_{D}^{(1)}$ was $2$-dimensional (it was the space of functions in $H_{D}$ with a certain value $a$ at $o$ and $w$ and a certain value $b$ at every $x\in D\setminus\{o,w\}$), in our case it is only $1$-dimensional: the two values $a$ and $b$ have to satisfy 2a+(s2)b=02a+(s-2)b=0.
The proof that $g$ preserves the three subspaces, that it preserves $Q$ on each of them, and that $H_{D}^{(3)}$ is orthogonal to both of the others works exactly as in the case $\alpha\neq-\frac{1}{s-1}$. The only part to change is where we check that $H_{D}^{(1)}$ and $H_{D}^{(2)}$ are orthogonal.
We use the same notation as before and fix arbitrary $h\in H_{D}^{(1)}$, $j\in H_{D}^{(2)}$. We let $a$ be the value of $h$ at $o$ and $w$, $b$ be the value of $h$ at any $x\in D\setminus\{o,w\}$ (they must satisfy 2a+(s2)b=02a+(s-2)b=0), and we assume without loss of generality $j(o)=1$ and j(w)=1j(w)=-1. We have